OpenStudy (carolina4567):

Write an equation of a line perpendicular to line AB below in slope-intercept form that passes through the point (7, 6). Line AB is shown. A is at 0, 1. B is at negative 2, 5. y = 0.5x + 2.5 y = 0.5x − 2.5 y = −2x + 20 y = 2x − 20

OpenStudy (carolina4567):

HELP

OpenStudy (carolina4567):

Anyone

OpenStudy (carolina4567):

@ganeshie8 @John_ES

OpenStudy (carolina4567):

I would think it is A what do u guys think

OpenStudy (carolina4567):

Please im really behind in gemoetry

OpenStudy (carolina4567):

@hartnn

OpenStudy (carolina4567):

@phi

OpenStudy (anonymous):

I think it might be A also... but i'm also really behind and kinda suck at geometry...

OpenStudy (carolina4567):

LOl we all suck at geometry

OpenStudy (carolina4567):

Jhon what do u think

OpenStudy (anonymous):

ikr >.< too bad the person who invented it is already dead because by now he'd have an angry mob after him

OpenStudy (john_es):

The correct answer is A, as @Rhea101 said.

OpenStudy (carolina4567):

Actually he'd be really dead by now

OpenStudy (carolina4567):

Thanks

OpenStudy (anonymous):

Lol true...

OpenStudy (anonymous):

Thanks for the medal

OpenStudy (carolina4567):

Got a few more if u dont mind jhon

OpenStudy (carolina4567):

Write an equation of a line perpendicular to y = 1 over 4x + 4 in slope-intercept form that passes through the point (−2, 6). y= 1 over 4x − 4 y= 1 over 4x + 6 y= −4x − 2 y= −4x + 1 over 4

OpenStudy (carolina4567):

Im thinking B

OpenStudy (john_es):

C, is the right answer.

OpenStudy (john_es):

When you put x=-2, you obtain y=6.

OpenStudy (phi):

they tell you A is (0, 1) B is (-2, 5) find the slope between the two points: change in y divided by change in x change in y is 1 - 5 = -4 change in x is 0 - (-2) = +2 change in y divided by change x = -4/2 = -2 the slope of a line perpendicular to a line with slope m is -1/m here, m = -2, and -1/m is -1/-2 = ½ or 0.5 that is the slope of your perpendicular line

OpenStudy (phi):

to find a line through through the point (7, 6) use y - y0= m(x- x0) where (x0,y0) is (7,6) and m (here) is 0.5 we get y - 6 = 0.5(x-7) y - 6 = 0.5x -3.5 y = 0.5x -3.5+6 and finally y = 0.5x +2.5